# Discrete Approximate Circle Bundles

### v2

@@ 3.3 Relationship To True Circle Bundles
In this section, we show that for sufficiently small $\alpha$, any discrete approximate local trivialization $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ corresponds to a true circle bundle. We build on previous results of [Tin22] and [SP23] which provide conditions under which approximate $O(2)$-cocycles and classifying maps can be uniquely identified with true ones.
The following construction and subsequent result, from [Tin22], describe a procedure for obtaining a classifying map from an approximate one:
Construction 3.13.
Given $A\in\mathbb{R}^{n\times n}$, compute the symmetrization $A_{s}=\frac{A+A^{T}}{2}$, then compute an orthogonal diagonalization $\Lambda D\Lambda^{T}$ of $A_{s}$ such that the eigenvalues appear in decreasing order along the diagonal of $D$. Finally, set $\Pi(A)=\Lambda J_{2}\Lambda^{T}$, where $J_{2}$ is the $n\times n$ matrix whose only non-zero entries are the first two diagonal entries, both equal to 1. Note that this construction naturally extends to the case $n=\infty$.
@@
Let $\Pi:\mathbb{R}^{\infty\times\infty}\to\text{Gr}(2)$ be the function described in Construction 3.13. For $\varepsilon\leq\frac{\sqrt{2}}{2}$, the restriction of $\Pi$ to $\text{Gr}(2)^{\varepsilon}$ is a retraction map. Furthermore, the induced map $\Pi_{*}:\text{Maps}(B,\text{Gr}(2)^{\varepsilon})\to\text{Maps}(B,\text{Gr}(2))$ descends to a bijection $\Pi_{*}:[B,\text{Gr}(2)^{\varepsilon}]\to[B,\text{Gr}(2)]$.
@@
Construction 3.15 defines a map $\widehat{\Pi}_{n}:S_{n}\to V(2,n)$ for each $n$, thereby inducing a map $\widehat{\Pi}:S\to V(2)$ (where $S$ is the direct limit of the sequence $S_{1}\subset S_{2}\subset\cdots$). Furthermore, $\widehat{\Pi}(P,A)$ is the closest matrix to $A$ in $\pi_{Gr}^{-1}(P)$ with respect to the Frobenius metric.
Finally, we have the following construction and subsequent result from [SP23]:
Construction 3.17.
Suppose $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$. Let $\{\Phi_{j}:U_{j}\to V(2)\}_{j\in J}$ be the maps defined by (15), and let $\Pi$ be as described in Construction 3.13. Define $f_{\Omega}=\Pi_{*}\circ\text{cl}(\Omega)$, then for each $j$, define $\widetilde{\Phi}_{j}:U_{j}\to V(2)$ by $\widetilde{\Phi}_{j}(b)=\hat{\Pi}(f_{\Omega}(b),\Phi_{j}(b))$. Finally, define $\widetilde{\Omega}\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ by $\widetilde{\Omega}_{jk}(b)=\widetilde{\Phi}_{j}^{T}(b)\widetilde{\Phi}_{k}(b)$.
@@
For $\varepsilon<\frac{1}{2}$, Construction 3.17 yields a map $\Pi_{Z}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\check{Z}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ such that the following diagram commutes:
@@
If $\varepsilon<\frac{\sqrt{2}}{2}$, any $\varepsilon$-approximate classifying map uniquely determines a true classifying map, hence can be identified with an isomorphism class of true circle bundles. Similarly, if $\varepsilon<\frac{1}{2}$, any $\varepsilon$-approximate $O(2)$ Cech 1-cocycle $\Omega$ uniquely determines a true circle bundle classified by $\Pi_{*}\circ\text{cl}(\Omega)$.
@@
Suppose $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ for $\alpha<\frac{\sqrt{2}}{20}$. Then any witness determines a true circle bundle, and the circle bundles determined by any two witnesses coincide. In this sense, $\{f_{j}\}$ uniquely determines a true circle bundle.
@@
Any witness $\Omega$ lies in $\check{Z}^{1}_{(3\sqrt{2})\alpha}(\mathcal{U};\underline{O(2)})$ by Proposition 3.4. Thus, if $\alpha<\frac{1}{6\sqrt{2}}$, then $(3\sqrt{2})\alpha<\frac{1}{2}$, so by Proposition 3.19 $\Omega$ determines a true circle bundle classified by $\Pi_{*}\circ\text{cl}(\Omega)$.
By Corollary 3.9, any two witnesses $\Omega,\Omega^{\prime}\in\check{Z}^{1}_{(3\sqrt{2})\alpha}(\mathcal{U};\underline{O(2)})$ satisfy $d_{Z}(\Omega,\Omega^{\prime})<2\varepsilon\sqrt{2}$. Thus, by Proposition 3.12, $\Pi_{*}\circ\text{cl}(\Omega)=\Pi_{*}\circ\text{cl}(\Omega^{\prime})$ as long as $\sqrt{2}(3\varepsilon\sqrt{2}+2\varepsilon\sqrt{2})=10\varepsilon\leq\frac{\sqrt{2}}{2}$, or equivalently, $\varepsilon\leq\frac{\sqrt{2}}{20}$.
@@
Suppose $\{f_{j}\},\{g_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ and $d_{T}(\{f_{j}\},\{g_{j}\})<\delta$. If $5\alpha+2\delta<\frac{\sqrt{2}}{4}$, then $\{f_{j}\}$ and $\{g_{j}\}$ both determine the same true circle bundle.
@@
By Proposition 3.10, any two witnesses $\Omega_{1},\Omega_{2}\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ that $\{f_{j}\},\{g_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ satisfy $d_{Z}(\Omega_{1},\Omega_{2})<2\sqrt{2}(\alpha+\delta)$. By Proposition 3.4, we have $\Omega_{1},\Omega_{2}\in\check{Z}^{1}_{(3\sqrt{2})\alpha}(\mathcal{U};\underline{O(2)})$, so Proposition 3.18 implies $\text{cl}(\Omega_{1})]$ and $\text{cl}(\Omega_{2})]$ coincide in $B,\text{Gr}(2)^{\sqrt{2}\varepsilon}]$, where $\varepsilon=(3\sqrt{2})\alpha+2\sqrt{2}(\alpha+\delta)$. Since $\Pi$ induces a bijection from $B,\text{Gr}(2)^{\sqrt{2}\varepsilon}]$ to $B,\text{Gr}(2)]$ by Proposition 3.19, $\Omega_{1}$ and $\Omega_{2}$ determine the same true circle bundle as long as
@@
Suppose $\{f_{j}\},\{g_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ are homotopic through $T_{\alpha}(\pi,\mathcal{U})$ for $\alpha<\frac{\sqrt{2}}{20}$. Then $\{f_{j}\}$ and $\{g_{j}\}$ determine the same true circle bundle.
@@
Suppose $\{h_{j}:\pi^{-1}(U_{j})\times I\to\mathbb{S}^{1}\}$ are continuous maps such that $\{h_{j}(\cdot,t)\}\in T_{\alpha}(\pi,\mathcal{U})$ for all $t$, and $h_{j}(x,0)=f_{j}(x)$ and $h_{j}(x,1)=g_{j}(x)$. Proposition 3.21 implies that for any $\alpha<\frac{\sqrt{2}}{20}$, there exists some $\delta>0$ such that $\{h_{j}(\cdot,t)\},\{h_{j}(\cdot,t+\Delta t)\}\in T_{\alpha}(\pi,\mathcal{U})$ determine the same true bundle as long as $d_{T}(\{h_{j}(\cdot,t)\},\{h_{j}(\cdot,t+\Delta t)\})<\delta$. By continuity, it follows that $\{h_{j}(\cdot,t)\}$ determines the same true bundle for all $t\in[0,1]$, and in particular that the true bundles determined by $\{f_{j}\}$ and $\{g_{j}\}$ coincide.
Next, we provide a construction for producing a family $\{\widetilde{f}_{j}\}\in T_{0}(\pi,\mathcal{U})$ from a family $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ when $\alpha$ is sufficiently small.
Construction 3.23.
Suppose $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$. Let $\Omega\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ be any witness, and let $\widetilde{\Omega}=\Pi_{Z}(\Omega)$. Choose a partition of unity $\{\rho_{j}\}_{j\in J}$ subordinate to $\mathcal{U}$, and for each $j$, define $\widetilde{f}_{j}:\pi^{-1}(U_{j})\to\mathbb{S}^{1}$ by
$\widetilde{f}_{j}(x)=\exp\left(\sum_{k\in J}\rho_{k}(\pi(x))\log\left(\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\right)\right)$ (17)
Note that if $\pi(x)\notin U_{j}\cap U_{k}$ then $\rho_{k}(\pi(x))=0$, so the maps are well-defined.
@@
Suppose $\Omega$ witnesses that $\{f_{j}\}_{j\in J}\in T_{\alpha}(\pi,\mathcal{U})$ and $\widetilde{\Omega}=\Pi_{Z}(\Omega)$. Then the maps $\{\widetilde{f}_{j}\}$ described in Construction 3.23 satisfy $\widetilde{f}_{j}(x)=\widetilde{\Omega}_{jk}(\pi(x))\widetilde{f}_{k}(x)$ for all $x\in\pi^{-1}(U_{j}\cap U_{k})$ and $U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$.
@@
Let $\widetilde{\Omega}=\Pi_{Z}(\Omega)$, and assume $\Omega$ decomposes as $\Omega_{jk}(b)=\exp(\Theta_{jk}(b))r(\omega_{jk})$ for some $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ and $\Theta\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{\mathbb{R}})$, where $\exp:\mathbb{R}\to SO(2)$ is as defined in (9) and $r:\mathbb{Z}_{2}\to O(2)$ is given by $r(\pm 1)=\left(\begin{array}[]{cc}1&0\\ 0&\pm 1\\ \end{array}\right)$. Then, for any $(jk)\in\mathcal{N}(\mathcal{U})$ and $x\in\pi^{-1}(U_{j}\cap U_{k})$, we have
@@
$=\quad\exp\left(\sum_{l}\rho_{l}(\pi(x))\log\left(\tilde{\Omega}_{jk}(b)\tilde{\Omega}_{kl}(b)f_{l}(x)\right)\right)$
@@
Suppose $\Omega$ witnesses that $\{f_{j}\}_{j\in J}\in T_{\alpha}(\pi,\mathcal{U})$ and $d_{Z}(\Omega,\widetilde{\Omega})<\delta$, where $\widetilde{\Omega}=\Pi_{Z}(\Omega)$. If $\delta+\alpha\sqrt{2}<2$, then for any partition of unity $\{\rho_{j}\}_{j\in J}$ subordinate to $\mathcal{U}$ and $x\in U_{j}$ for some $j\in J$, $\widetilde{f}_{j}(x)$ as defined by (17) is equal to the unique weighted Karcher mean of $\{\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\}_{k\in K_{x}}$ with weights $\{\rho_{k}(\pi(x))\}_{k\in K_{x}}$, where $K_{x}=\{k\in J:x\in U_{j}\cap U_{k}\}$. Furthermore $d_{T}(\{f_{j}\},\{\widetilde{f}_{j}\})<\alpha+\frac{\delta}{\sqrt{2}}$, and the maps $\{\widetilde{f}_{j}\}_{j\in J}$ obtained by different choices of partition of unity are homotopic through $T_{0}(\pi,\mathcal{U})$.
@@
For any $(jk)\in\mathcal{N}(\mathcal{U})$ and $x\in\pi^{-1}(U_{j}\cap U_{k})$, we have
$|f_{j}(x)-\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)|$ $\leq|f_{j}(x)-\Omega_{jk}f_{k}(x)|+|\Omega_{jk}f_{k}(x)-\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)|$
$<\alpha+\frac{|\Omega_{jk}-\widetilde{\Omega}_{jk}(\pi(x))|_{F}}{\sqrt{2}}$
$\leq\alpha+\frac{\delta}{\sqrt{2}}$
Thus, by Proposition A.11, $\widetilde{f}_{j}(x)$ as defined in (17) is the unique weighted Karcher mean of the set $\{\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\}_{k\in K_{x}}$ as long as $\alpha+\frac{\delta}{\sqrt{2}}<\sqrt{2}$. Clearly we have $d_{T}(\{f_{j}\}_{j\in J},\{\widetilde{f}_{j}\}_{j\in J})\leq\alpha+\frac{\delta}{\sqrt{2}}$ in this case.
Now, suppose $\{\rho^{\prime}_{j}\}_{j\in J}$ is another partition of unity subordinate to $\mathcal{U}$. Each partition of unity can be viewed as the component functions of a map $\rho,\rho^{\prime}:B\to\Delta^{n-1}$. Since $\Delta^{n-1}$ is convex, we can form the straight line homotopy $h_{t}:B\to\Delta^{n-1}$ from $\rho$ to $\rho^{\prime}$. Replacing the components of $\rho$ with those of $h_{t}$ in (17) for each $t\in[0,1]$ gives a continuously varying family in $T_{0}(\pi,\mathcal{U})$, so $\{f_{j}\}$ and $\{\widetilde{f}_{j}\}$ are homotopic through $T_{0}(\pi,\mathcal{U})$.
@@
For $\alpha<\frac{\sqrt{2}}{28}$, Construction 3.23 defines a map $\Pi_{T}:T_{\alpha}(\pi,\mathcal{U})\to T_{0}(\pi,\mathcal{U})$ such that for any $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$, the families $\{f_{j}\}$ and $\Pi_{T}(\{f_{j}\})$ determine the same true bundle.

### v3

@@ 3.5 Relationship To True Circle Bundles
In this section, we show that for sufficiently small $\alpha>0$, any system of discrete approximate local angle functions $\mathfrak{F}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ uniquely corresponds to a true circle bundle. From this we deduce conditions under which a discrete approximate circle bundle can be identified with an isomorphism class of classical circle bundles. We build on previous results of [30] and [24] which provide conditions under which approximate $O(2)$-cocycles and classifying maps can be uniquely identified with true ones.
The following construction and subsequent result [30, Lemma 1] describe a procedure for obtaining a classifying map from an approximate one:
Construction 3.24.
Given $A\in\mathbb{R}^{n\times n}$, compute its symmetrization $A_{s}=\frac{A+A^{T}}{2}$, and an orthogonal diagonalization $A_{s}=\Lambda D\Lambda^{T}$ such that the eigenvalues of $A_{s}$ appear in decreasing order along the diagonal of $D$. Finally, set $\Pi(A)=\Lambda J_{2}\Lambda^{T}$, where $J_{2}$ is the $n\times n$ matrix whose only non-zero entries are the first two diagonal entries, both equal to 1. Note that this construction naturally extends to the case $n=\infty$.
@@
For $\varepsilon\leq\frac{\sqrt{2}}{2}$, Construction 3.24 yields a deformation retraction $\Pi:\text{Gr}(2)^{\varepsilon}\to\text{Gr}(2)$. Hence, the induced map $\Pi_{*}:\text{Maps}(B,\text{Gr}(2)^{\varepsilon})\to\text{Maps}(B,\text{Gr}(2))$ descends to a natural bijection $\Pi_{*}:[B,\text{Gr}(2)^{\varepsilon}]\to[B,\text{Gr}(2)]$.
@@
Construction 3.26 defines a map $\widehat{\Pi}_{n}:S_{n}\to V(2,n)$ for each $n$, thereby inducing a map $\widehat{\Pi}:S\to V(2)$ (where $S$ is the direct limit of the sequence $S_{1}\subset S_{2}\subset\cdots$). Furthermore, $\widehat{\Pi}(P,A)$ is the point in $\pi_{Gr}^{-1}(P)$ closest to $A$ with respect to the Frobenius metric.
Finally, we have the following construction and subsequent result [24, Theorem 4.27]:
Construction 3.28.
Suppose $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for $\varepsilon<\frac{1}{2}$. Let $\{\Psi_{j}:U_{j}\to V(2)\}_{j\in\mathbb{N}}$ be the maps defined by (21), and let $\Pi$ be as described in Construction 3.24. Define $f_{\Omega}=\Pi_{*}\circ\text{cl}(\Omega)$, and for each $j$ define $\widetilde{\Psi}_{j}:U_{j}\to V(2)$ by $\widetilde{\Psi}_{j}(b)=\hat{\Pi}(f_{\Omega}(b),\Psi_{j}(b))$. Finally, define $\widetilde{\Omega}\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ by $\widetilde{\Omega}_{jk}(b)=\widetilde{\Psi}_{j}(b)^{T}\widetilde{\Psi}_{k}(b)$.
@@
For $\varepsilon<\frac{1}{2}$, Construction 3.28 yields a map $\Pi_{Z}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\check{Z}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ such that the following diagram commutes:
@@
If $\varepsilon<\frac{\sqrt{2}}{2}$, any $\varepsilon$-approximate classifying map uniquely determines a true classifying map, hence can be identified with an isomorphism class of true circle bundles. Similarly, if $\varepsilon<\frac{1}{2}$, any $\varepsilon$-approximate $O(2)$ Čech 1-cocycle $\Omega$ uniquely determines a true circle bundle classified by $\Pi_{*}\circ\text{cl}(\Omega)$.
@@
Suppose $\mathfrak{F}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for $\alpha<\frac{\sqrt{2}}{20}$. Then any witness determines a true circle bundle, and the circle bundles determined by any two witnesses coincide. In this sense, $\mathfrak{F}$ uniquely determines a true circle bundle.
@@
Any witness $\Omega$ lies in $\check{Z}^{1}_{(3\sqrt{2})\alpha}\left(\mathcal{U};\underline{O(2)}\right)$ by Proposition 3.15. Thus, if $\alpha<\frac{1}{6\sqrt{2}}$, then $(3\sqrt{2})\alpha<\frac{1}{2}$, so by Proposition 3.30 $\Omega$ determines a true circle bundle classified by $\Pi_{*}\circ\text{cl}(\Omega)$.
By Corollary 3.20, any two witnesses $\Omega,\Omega^{\prime}\in\check{Z}^{1}_{(3\sqrt{2})\alpha}\left(\mathcal{U};\underline{O(2)}\right)$ satisfy $d_{Z}(\Omega,\Omega^{\prime})<2\varepsilon\sqrt{2}$. Thus, by Proposition 3.23, $\Pi_{*}\circ\text{cl}(\Omega)=\Pi_{*}\circ\text{cl}(\Omega^{\prime})$ as long as $\sqrt{2}(3\varepsilon\sqrt{2}+2\varepsilon\sqrt{2})=10\varepsilon\leq\frac{\sqrt{2}}{2}$, or equivalently, $\varepsilon\leq\frac{\sqrt{2}}{20}$.
@@
Suppose $\mathfrak{F},\mathfrak{G}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ and $d_{\Gamma}(\mathfrak{F},\mathfrak{G})<\delta$. If $5\alpha+2\delta<\frac{\sqrt{2}}{4}$, then $\mathfrak{F}$ and $\mathfrak{G}$ both determine the same true circle bundle.
@@
By Corollary 3.21, any two witnesses $\Omega_{1},\Omega_{2}\in\check{C}^{1}\left(\mathcal{U};\underline{O(2)}\right)$ for $\mathfrak{F},\mathfrak{G}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ satisfy $d_{Z}(\Omega_{1},\Omega_{2})<2\sqrt{2}(\alpha+\delta)$. By Proposition 3.15, we have $\Omega_{1},\Omega_{2}\in\check{Z}^{1}_{(3\sqrt{2})\alpha}\left(\mathcal{U};\underline{O(2)}\right)$, so Proposition 3.29 implies $\text{cl}(\Omega_{1})]$ and $\text{cl}(\Omega_{2})]$ coincide in $\left[B,\text{Gr}(2)^{\sqrt{2}\varepsilon}\right]$, where $\varepsilon=(3\sqrt{2})\alpha+2\sqrt{2}(\alpha+\delta)$. Since $\Pi$ induces a bijection from $\left[B,\text{Gr}(2)^{\sqrt{2}\varepsilon}\right]$ to $B,\text{Gr}(2)]$ by Proposition 3.30, $\Omega_{1}$ and $\Omega_{2}$ determine the same true circle bundle as long as
@@
Suppose $\mathfrak{F}=\{f_{j}\}_{j\in J},\mathfrak{G}=\{g_{j}\}_{j\in J}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for some $\alpha<\frac{\sqrt{2}}{20}$, and that there are homotopies $\{h_{j}:U_{j}\times I\to\mathbb{S}^{1}\}_{j\in J}$ such that $h_{j}(\cdot,0)=f_{j}$, $h_{1}(\cdot,1)=g_{j}$ and $\{h_{j}(\cdot,t)\}_{j\in J}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for all $t\in I$. Then $\mathfrak{F}$ and $\mathfrak{G}$ determine the same true circle bundle.
@@
Suppose $\{h_{j}:\pi^{-1}(U_{j})\times I\to\mathbb{S}^{1}\}_{j\in J}$ are continuous maps such that $\{h_{j}(\cdot,t)\}_{j\in J}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for all $t$, and $h_{j}(x,0)=f_{j}(x)$ and $h_{j}(x,1)=g_{j}(x)$. Proposition 3.32 implies that for any $\alpha<\frac{\sqrt{2}}{20}$, there exists some $\delta>0$ such that for any $\Delta t>0$, $\{h_{j}(\cdot,t)\}_{j\in J},\{h_{j}(\cdot,t+\Delta t)\}_{j\in J}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ determine the same true bundle as long as $d_{\Gamma}(\{h_{j}(\cdot,t)\}_{j\in J},\{h_{j}(\cdot,t+\Delta t)\}_{j\in J})<\delta$. By continuity, it follows that $\{h_{j}(\cdot,t)\}_{j\in J}$ determines the same true bundle for all $t\in[0,1]$, and in particular that the true bundles determined by $\{f_{j}\}_{j\in J}$ and $\{g_{j}\}_{j\in J}$ coincide.
Remark 3.34.
Discrete approximate local angle functions satisfying the hypotheses of Corollary 3.33 above for some $\alpha$ are said to be homotopic through $\Gamma_{\alpha}(\pi,\mathcal{U})$.
Given a function $\pi:X\to B$ between metric spaces and an open cover $\mathcal{U}=\{U_{j}\}$ of $B$, we define $\Gamma_{0}(\pi,\mathcal{U})$ to be the set of families $\{f_{j}:U_{j}\to\mathbb{S}^{1}\}_{j\in J}$ for which there exists some Čech cocycle $\Omega\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ satisfying $f_{j}(b)=\Omega_{jk}(b)f_{k}(b)$ for all $U_{jk}\in\mathcal{N}(\mathcal{U})$ and $b\in U_{jk}$. Note that $\Omega$ need not be locally constant. We provide a construction for producing a family $\{\widetilde{f}_{j}\}_{j\in J}\in\Gamma_{0}(\pi,\mathcal{U})$ from a family $\{f_{j}\}_{j\in J}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ when $\alpha$ is sufficiently small and $\mathcal{U}=\{U_{j}\}_{j\in J}$ is locally finite. Our construction uses the notion of a Frechet mean [13] for angular values in $\mathbb{S}^{1}$, which generalizes the notion of a centroid in Euclidean space:
Definition 3.35.
Let $M$ be a complete metric space, and let $m_{1},...,m_{N}\in M$. For any $p\in M$, the (weighted) Frechet variance of the set $\{m_{1},...,m_{N}\}$ at $p$ with weights $\{w_{1},...,w_{N}\}$ is defined by
$F_{V}(p)=\sum_{n=1}^{N}w_{n}d_{M}(p,x_{n})^{2}$
The point $p$ is called a (weighted) Karcher mean of $\{m_{1},...,m_{N}\}$ if it minimizes $F_{V}$ on $M$. If $p$ is the unique weighted Karcher mean of $\{m_{1},...,m_{N}\}$, it is called the (weighted) Frechet mean of $\{m_{1},...,m_{N}\}$.
Construction 3.36.
Suppose $\{f_{j}\}_{j\in J}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for some locally-finite good open cover $\mathcal{U}$. Let $\Omega\in\check{C}^{1}\left(\mathcal{U};\underline{O(2)}\right)$ be any witness, and let $\widetilde{\Omega}=\Pi_{Z}(\Omega)$. Choose a partition of unity $\{\rho_{j}\}_{j\in J}$ subordinate to $\mathcal{U}$, and for each $j$, define $\widetilde{f}_{j}:\pi^{-1}(U_{j})\to\mathbb{S}^{1}$ by
$\widetilde{f}_{j}(x)\quad=\quad\exp\left(\sum_{k\in J}\rho_{k}(\pi(x))\log\left(\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\right)\right)$ (23)
Here we fix a branch of log for each $x\in X$ and each $j\in J$ as follows. Since $\mathcal{U}$ is locally finite, then the set $\left\{\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\in\mathbb{S}^{1}\,:\,k\in J\right\}$ is finite, and thus is contained in the domain of some branch of log. Note that if $\pi(x)\notin U_{jk}$ then $\rho_{k}(\pi(x))=0$, so $\tilde{f}_{j}(x)$ is well-defined.
@@
Suppose $\Omega$ witnesses that $\mathfrak{F}=\{f_{j}\}_{j\in J}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for some good open cover $\mathcal{U}$. Then the functions $\{\widetilde{f}_{j}\}_{j\in J}$ described in Construction 3.36 form a system $\widetilde{\mathfrak{F}}\in\Gamma_{0}(\pi,\mathcal{U})$.
@@
Let $\exp:\mathbb{R}\to SO(2)$ be as defined in (8) and define $r:\mathbb{Z}_{2}\to O(2)$ by $r(\pm 1)=\left(\begin{array}[]{cc}1&0\\ 0&\pm 1\\ \end{array}\right)$. Let $\widetilde{\Omega}=\Pi_{Z}(\Omega)$. Define $\omega=\det_{*}(\Omega)$, and for each $U_{jk}\in\mathcal{N}(\mathcal{U})$, let $R_{jk}:U_{jk}\to SO(2)$ be defined by $R_{jk}(b)=\Omega_{jk}(b)r(\omega_{jk})$. Since $\mathcal{U}$ is a good open cover, each intersection $U_{jk}$ is contractible, so each $R_{jk}$ has a continuous lift $\Theta_{jk}:U_{jk}\to\mathbb{R}$ satisfying $\exp(\Theta_{jk}(b))=R_{jk}(b)$. Thus, $\Omega$ decomposes as $\Omega_{jk}(b)=\exp(\Theta_{jk}(b))r(\omega_{jk})$. For each $x\in X$ and $j\in J$ such that $\pi(x)\in U_{j}$, choose a branch of $\log$ whose domain contains the (finite) set $\{\tilde{\Omega}_{jk}(\pi(x))f_{k}(x)\in\mathbb{S}^{1}:\pi(x)\in U_{k}\}$. Note that whenever $\pi(x)=b\in\pi^{-1}(U_{j}\cap U_{k})$ for some $j,k\in J$, we have
$\log\left(\widetilde{\Omega}_{jk}(b)f_{k}(x)\right)\ =\ \kappa_{jk}(b)+\Theta_{jk}(b)+\omega_{jk}\log\left(f_{k}(x)\right)$
for some $\kappa_{jk}(\pi(x))\in\mathbb{Z}=\ker(\exp)$, so
@@
$=\quad\exp\left(\sum_{l}\rho_{l}(b)\log\left(\tilde{\Omega}_{jk}(b)\tilde{\Omega}_{kl}(b)f_{l}(x)\right)\right)$
@@
Suppose $\Omega$ witnesses that $\{f_{j}\}_{j\in J}\in T_{\alpha}(\pi,\mathcal{U})$ and $d_{Z}(\Omega,\widetilde{\Omega})<\delta$, where $\widetilde{\Omega}=\Pi_{Z}(\Omega)$. If $\delta+\alpha<\sqrt{2}$, then for any partition of unity $\{\rho_{j}\}_{j\in J}$ subordinate to $\mathcal{U}$ and $x\in U_{j}$ for some $j\in J$, $\widetilde{f}_{j}(x)$ as defined by (23) is equal to the Frechet mean of $\{\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\}_{k\in K_{x}}$ with weights $\{\rho_{k}(\pi(x))\}_{k\in K_{x}}$, where $K_{x}=\{k\in J:x\in U_{jk}\}$. Furthermore $d_{\Gamma}(\{f_{j}\}_{j\in J},\{\widetilde{f}_{j}\}_{j\in J})<\alpha+\delta$, and the functions $\{\widetilde{f}_{j}\}_{j\in J}$ obtained by different choices of partition of unity are homotopic through $\Gamma_{0}(\pi,\mathcal{U})$.
@@
For any $(jk)\in\mathcal{N}(\mathcal{U})$ and $x\in\pi^{-1}(U_{jk})$, we have
$|f_{j}(x)-\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)|\quad$ $\leq\quad|f_{j}(x)-\Omega_{jk}f_{k}(x)|+|\Omega_{jk}f_{k}(x)-\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)|$
$<\quad\alpha+\|\Omega_{jk}-\widetilde{\Omega}_{jk}(\pi(x))\|_{F}$
$\leq\quad\alpha+\delta$
Thus, by Proposition A.13, $\widetilde{f}_{j}(x)$ as defined in (23) is the unique weighted Karcher mean of the set $\{\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\}_{k\in K_{x}}$ as long as $\alpha+\delta<\sqrt{2}$. Clearly we have $d_{\Gamma}(\{f_{j}\}_{j\in J},\{\widetilde{f}_{j}\}_{j\in J})\leq\alpha+\delta$ in this case.
Now, suppose $\{\rho^{\prime}_{j}\}_{j\in J}$ is another partition of unity subordinate to $\mathcal{U}$. Each partition of unity can be viewed as the component functions of maps $\rho,\rho^{\prime}:B\to\Delta^{J}$ to the full simplex on $J$. Since $\Delta^{J}$ is convex, we can form the straight line homotopy $h_{t}:B\to\Delta^{J}$ from $\rho$ to $\rho^{\prime}$. Replacing the components of $\rho$ with those of $h_{t}$ in (23) for each $t\in[0,1]$ gives a continuously varying family in $\Gamma_{0}(\pi,\mathcal{U})$, so $\{f_{j}\}_{j\in J}$ and $\{\widetilde{f}_{j}\}_{j\in J}$ are homotopic through $\Gamma_{0}(\pi,\mathcal{U})$.
@@
For $\alpha<\frac{1}{28}$, Construction 3.36 defines a function $\Pi_{\Gamma}:\Gamma_{\alpha}(\pi,\mathcal{U})\to\Gamma_{0}(\pi,\mathcal{U})$ such that for any $\mathfrak{F}\in\Gamma_{\alpha}(\pi,\mathcal{U})$, the systems $\mathfrak{F}$ and $\Pi_{\Gamma}(\mathfrak{F})$ determine the same true bundle.